# Statistical Finite Elements via Interacting Particle Langevin Dynamics

### v1

@@ 1 Introduction
where $k$ is a (regular enough) kernel function (see section 2.1 for details). A data generating process can then be defined through the continuous observation operator $\mathcal{H}\colon\Omega\rightarrow\mathbb{R}^{n_{y}}$
@@
The equations (1)–(2) specify a statistical model, which can be used to infer the latent PDE solution $u$. This is a standard problem, having been the subject of much recent attention with, for example, the statFEM [4, 16, 24]. While the problem of sampling (or inferring) the statFEM solution is relatively well understood, the joint estimation of any PDE parameters and the model solution remains an open problem. Denote by $\mathbf{z}$ these unknown parameters, which may comprise any unknown model components, e.g. PDE parameters $\theta$, or the forcing function $f$. Discretising (1) using statFEM, we obtain a model state vector $\mathbf{u}$, and conditional distribution $p(\mathbf{u}|\mathbf{z})$. To estimate the parameters of the PDE (that are denoted as $\mathbf{z}$), we aim at maximising the marginal likelihood by solving the following optimisation problem
@@
which is termed the maximum marginal likelihood estimation (MMLE) problem. When we include a prior distribution in the model, the marginal maximum a posteriori (MMAP) problem can be solved by maximising the marginal posterior distribution
$\mathbf{z}_{\mathrm{MMAP}}=\operatorname*{\arg\!\max}_{\mathbf{z}\in\mathcal{Z}}\log p(\mathbf{z}|\mathbf{y}),\quad\text{where}\quad p(\mathbf{z}|\mathbf{y}):=\int p(\mathbf{y}|\mathbf{u})\,p(\mathbf{u}|\mathbf{z})p(\mathbf{z})\mathrm{d}\mathbf{u}.$ (4)
@@ 1.2 Marginal MLE and MAP
The problem of joint state and parameter estimation has received attention from various communities, including physics [1], chemistry [15], electrical engineering [39, 40], data assimilation [30, 21], and statistics [27]. For the extension to PDE systems, Bayesian inverse problems (BIP) for PDEs tackle the problem of parameter estimation from data or inversion, but typically assume a known (and deterministic) PDE forward model, which is restrictive in the modelling of uncertainties. By constructing a probabilistic forward model based on our assumed PDE, we allow for principled accounting for model misspecification, and joint estimation of parameters and uncertainty quantification (UQ) for our PDE solution.
@@ 1.3 Our contribution
  • We develop general algorithms for forcing estimation in linear Poisson PDE (section 3.1) and nonlinear Poisson PDE (section 3.2). Our choice of Poisson PDEs is for demonstration since our algorithms can be similarly adapted to other PDEs. We develop preconditioning strategies within these sections to improve the condition number of the problem as well as proposing several Gaussian approximations for the nonlinear case.
  • We then develop algorithms for solving inverse problems (section 4), i.e., estimating the parameter of the forward map given partial observations of the PDE solution for the linear case. We introduce warm-start strategies to avoid the numerical difficulties caused by the ill-conditioning of the problem.
@@
  • We provide experiments to demonstrate the order of convergence proved in Theorem 1 using an analytical example, as well as the benefit preconditioning through computation of the condition number (section 6.1). We compare the different nonlinear approximations for forcing estimation in section 6.2, and show the benefit of using a warm-start for solving the inverse problem in section 6.3.
This paper is organised as follows: in section 2 we introduce the statistical finite element method for linear and nonlinear PDEs before introducing Langevin dynamics for statFEM and the interacting particle Langevin algorithm. We then show how the statFEM construction can be used to embed the probabilistic model for the PDE within the IPLA for posterior sampling. Section 3 derives the general algorithm for forcing estimation, which is developed for both linear and nonlinear Poisson problems in sections 3.1 and 3.2, respectively. In section 4 we derive the generic algorithm this time for solving the inverse problem of PDE parameter estimation, then apply this to linear Poisson diffusivity estimation in 4.1. Theoretical results are derived in section 5 and experimental results for forcing estimation and the inverse problem are given in section 6. We conclude the paper in section 7.
@@ 2.1 Statistical Finite Elements for Linear PDEs
$\mathcal{A}_{\theta}(u_{h},\phi_{j})=\sum_{i=1}^{n_{u}}\hat{u}_{i}\mathcal{A}_{\theta}(\phi_{i},\phi_{j})=\left\langle f,\phi_{j}\right\rangle+\left\langle\epsilon,\phi_{j}\right\rangle,\quad\forall j\in\{1,\dots,n_{u}\}.$ (8)
@@
For setting the homogeneous Dirichlet boundary conditions, we set the rows corresponding to boundary nodes to the row of the identity matrix, and the forcing vector to zero. This is easily adapted for non-zero boundary conditions by setting the forcing vector value to the non-zero boundary value. We choose to constrain the covariance function $k$ to be zero on the boundary, so that boundary conditions are imposed exactly. To achieve this, we form an approximation to the GP, as seen in [36], which has the benefit in our case of being constrained to zero covariance on the boundary. The eigenfunctions of the Laplacian operator (essentially the vibrational modes of the domain) form the basis, which are weighted based on the spectrum of the covariance function. This approximation is given by
@@
where the function $g_{l}(x)=\sum_{l=1}^{n_{u}}\tilde{g}_{li}\phi_{i}(x)$ is the $L^{2}$-normalised FEM approximation of the $l^{\text{th}}$ eigenfunction of the Laplacian operator, $\lambda_{l}$ the corresponding eigenvalue, and $m$ the rank of the approximation. The weighting function $S(\cdot)$ is the spectral density of the covariance function $k$. Note this requires an isotropic covariance function, i.e. such that we can write $k(x,x^{\prime})=k(\|x-x^{\prime}\|)$. This is detailed in Appendix E.
2.2 Statistical Finite Elements for Nonlinear PDEs
As we have seen in section 2.1, if the PDE is defined by a linear differential operator the prior distribution induced by the GP noise is also a Gaussian. This is not the case in general. A nonlinear differential operator will induce a (typically non-analytic) non-Gaussian prior over the FEM coefficients. A stochastic PDE with the nonlinear differential operator $\mathcal{F}_{\theta}$ and forcing $f$
$\mathcal{F}_{\theta}(u)=f+\epsilon,\quad\epsilon\sim\mathcal{GP}(0,k(\cdot,\cdot)),$ (12)
can be discretised via the statFEM method giving a system of nonlinear algebraic equations $\mathbf{F}_{\theta}(\mathbf{u})=\mathbf{b}+\mathbf{e},\quad\mathbf{e}\sim\mathcal{N}(\mathbf{0},\mathbf{G})$ (see Appendix C for deviation). The induced prior distribution can be found via a transformation of variables [6], from distribution $\mathbf{E}\sim f_{\mathbf{E}}(\mathbf{e})$ to $\mathbf{U}\sim f_{\mathbf{U}}(\mathbf{u})$ where $\mathbf{u}=g(\mathbf{e})$,
Using this transformation of variables with $f_{\mathbf{E}}=\mathcal{N}(\mathbf{0},\mathbf{G})$ and $g^{-1}(\mathbf{u})=\mathbf{F}_{\theta}(\mathbf{u})-\mathbf{b}$, we can see how the resulting density of the solution depends on the nonlinear differential operator and its Jacobian:
$p(\mathbf{u}|\theta,\mathbf{b})=\det|\mathbf{J}_{\theta}(\mathbf{u})|\frac{1}{Z}\exp\left\{-\frac{1}{2}\left(\mathbf{F}_{\theta}(\mathbf{u})-\mathbf{b}\right)^{\top}\mathbf{G}^{-1}\left(\mathbf{F}_{\theta}(\mathbf{u})-\mathbf{b}\right)\right\},$ (14)
where the constant $Z=(2\pi)^{-n_{u}/2}\det|\mathbf{G}|^{-1/2}$, and the Jacobian of the nonlinear transform $\mathbf{F}_{\theta}$ is denoted $\mathbf{J}_{\theta}(\mathbf{u})=\frac{\partial\mathbf{F}_{\theta}}{\partial\mathbf{u}}\big{|}_{\mathbf{u}}$. This prior, in general, is non-Gaussian due to the Jacobian determinant term.
@@ 2.3 Statistical Finite Elements via Langevin Dynamics
where $\left\{\bm{\zeta}_{k}\right\}_{k\in\mathbb{N}}$ is a sequence of i.i.d. standard Gaussian random variables, and $\gamma>0$ controls the degree of time-discretisation. This discretisation induces a bias, the degree of which is determined by the step-size $\gamma$. For some applications this bias may be unacceptable, and can be corrected by a Metropolis-Hastings style accept/reject step for each iteration of the algorithm (this is known as Metropolis adjusted Langevin algorithm (MALA) [23]) under the condition that the unnormalised density can be evaluated directly (not just the score as for ULA). ULA does not perform this accept/reject stage and is therefore more computationally efficient than MALA, at the cost of introducing discretization bias [10, 11, 18].
@@
Preconditioning can improve the stability and convergence of the sampling scheme, whilst leaving the stationary measure invariant [4]. Section 3 shows how preconditioning is applied.
@@ 2.4 Interacting Particle Langevin Dynamics
Whilst sections below will detail the IPLA algorithm for forcing estimation and solving the inverse problem, we provide a brief overview here. IPLA simultaneously evolves a set of particles with Langevin dynamics, which are used to estimate the gradient of the marginal likelihood with respect to parameters. A system of $N$ particles, $\{\mathbf{u}^{(n)}\}_{n=1}^{N}$ are simulated from the Langevin dynamics described in (18), with independent Brownian motions $(\mathbf{B}_{t}^{(n)})_{t\geq 0}$ for $n=1,\ldots,N$. These particles and parameters are jointly performing gradient flow on the negative joint log likelihood where the parameter process is driven with a scaled Brownian motion $\sqrt{{2}/{N}}\mathrm{d}\mathbf{B}_{t}^{(0)}$. This noise term, which is the main difference from the methods introduced in [29], allows for non-asymptotic analysis, and in [3], it is shown that given Lipschitz and strong convexity assumptions on the potential, the resulting parameter estimate will converge to the true MMLE, and the rate of convergence is derived in terms of the Lipschitz and strong convexity constants, and time-discretization step-size. The method has also shown to be adaptable to the nondifferentiable and underdamped cases, see, e.g. [19, 31]. Sections 3 and 4 give the resulting interacting particle systems for estimation of the forcing function and solving the inverse problem, respectively.
3 Estimating Forcing via Interacting Particle Langevin Dynamics
In this section, we focus on estimating the forcing function $f$ for fixed $\theta$. We place a prior on the finite-element coefficients $\mathbf{b}$ that form the forcing function basis expansion $f(x)\approx\sum_{i=1}^{n_{u}}[\mathbf{b}]_{i}\phi_{i}(x)$. Given $\theta$ is fixed, we introduce a simplified notation:
$\Phi(\mathbf{u},\mathbf{b}):=-\log p(\mathbf{u}|\theta,\mathbf{b})-\log p(\mathbf{b}),\quad\text{and}\quad\Phi^{y}(\mathbf{u},\mathbf{b}):=\Phi(\mathbf{u},\mathbf{b})-\log p(\mathbf{y}|\mathbf{u}),$ (20)
where $\Phi(\mathbf{u},\mathbf{b}),\Phi^{y}(\mathbf{u},\mathbf{b})$ are the prior and posterior potentials respectively. We can now write the general form of the algorithm for forcing estimation using the posterior potential:
$\mathbf{b}_{k+1}$ $=\mathbf{b}_{k}-\frac{\gamma}{N}\sum_{n=1}^{N}\nabla_{\mathbf{b}}\Phi^{y}(\mathbf{u}_{k}^{(n)},\mathbf{b}_{k})+\sqrt{\frac{2\gamma}{N}}\,\bm{\zeta}_{k+1}^{(0)}$ (21)

### v2

@@ 1 Introduction
where $k$ is a (regular enough) kernel function (see section 2.1 for details). A data generating process can then be defined through the continuous observation operator $\mathcal{H}\colon L^{2}(\Omega)\rightarrow\mathbb{R}^{n_{y}}$
@@
The equations (1)–(2) specify a statistical model, which can be used to infer the latent PDE solution $u$. This is a standard problem, having been the subject of much recent attention with, for example, the statFEM [4, 18, 28]. While these methods can be generically used for sampling the parameters jointly, there is a gap in the literature for dedicated methods to obtain point estimates of the parameters directly, which may avoid a separate procedure to get maximisers from samples. Denote by $\mathbf{z}$ these unknown parameters, which may comprise any unknown model components, e.g. PDE parameters $\theta$, or the forcing function $f$. Discretising (1) using statFEM, we obtain a model state vector $\mathbf{u}$, and conditional distribution $p(\mathbf{u}|\mathbf{z})$. To estimate the parameters of the PDE (that are denoted as $\mathbf{z}$), we aim at maximising the marginal likelihood by solving the following optimisation problem
@@
which is termed the marginal maximum likelihood estimation (MMLE) problem. When we include a prior distribution in the model, the marginal maximum a posteriori (MMAP) problem can be solved by maximising the marginal posterior distribution
$\mathbf{z}_{\mathrm{MMAP}}=\operatorname*{\arg\!\max}_{\mathbf{z}\in\mathcal{Z}}\log p(\mathbf{z},\mathbf{y}),\quad\text{where}\quad p(\mathbf{z},\mathbf{y}):=\int p(\mathbf{y}|\mathbf{u})\,p(\mathbf{u}|\mathbf{z})p(\mathbf{z})\mathrm{d}\mathbf{u}.$ (4)
@@ 1.2 Marginal MLE and MAP
The problem of joint state and parameter estimation has received attention from various communities, including physics [1], chemistry [17], electrical engineering [47, 48], data assimilation [36, 25], and statistics [32]. For the extension to PDE systems, Bayesian inverse problems for PDEs tackle the problem of parameter estimation (i.e. $\mathbf{z}=\theta$) from data or inversion, but typically assume a known (and deterministic) PDE forward model, which is restrictive in the modelling of uncertainties. Estimation of PDE forcing/source terms from partial observations (i.e. $\mathbf{z}=f$) is relevant to climate modelling, for example the estimation of greenhouse gas sources from satellite observations [37] or identification of pollutant sources [6]. This is also of interest to the engineering community, for example in structural health monitoring for estimating unknown forcing from sensor measurements [22]. By constructing a probabilistic forward model based on our assumed PDE, we allow for principled accounting for model misspecification, and joint estimation of parameters and uncertainty quantification for our PDE solution.
@@ 1.3 Our contribution
  • We develop an algorithm for forcing estimation (section 3.1) and solving the inverse problem (section 3.2) in section 3, which is applied to the linear Poisson PDE. Our choice of Poisson PDEs is for demonstration since our algorithms can be similarly adapted to other PDEs.
  • We develop preconditioning strategies within this section to improve the condition number of the problem. We also introduce warm-start strategies to avoid the numerical difficulties caused by the ill-conditioning of the problem.
@@
  • We provide experiments to demonstrate the order of convergence proved in Theorem 1 using an analytical example, as well as the benefit preconditioning through computation of the condition number (section 5.1). We show the benefit of using a warm-start for solving the inverse problem in section 5.2.
  • We develop an algorithm for forcing estimation in the nonlinear statFEM setting (section 6.1), which we apply to the nonlinear Poisson PDE (section 6.2). We compare three different nonlinear approximations for forcing estimation in section 6.3.
This paper is organised as follows: in section 2 we introduce the statistical finite element method for linear PDEs (section 2.1) before introducing Langevin dynamics for statFEM (section 2.2) and the interacting particle Langevin algorithm (section 2.3). Section 3 shows how the statFEM construction can be used to embed the probabilistic model for the linear Poisson PDE within the IPLA for MMLE/MMAP estimation. The algorithms and preconditioners are given in section 3.1 for forcing estimation, and section 3.2 for solving the inverse problem. Theoretical results are derived in section 4. Experimental results for the linear Poisson PDE are given in section 5, including the forcing estimation (section 5.1), and the inverse problem (section 5.2) results. In section 6, we extend the methodology to nonlinear statFEM. Section 6.1 derives the form of the statFEM prior for the nonlinear Poisson PDE, and section 6.2 outlines three different approximations to this prior for use in the IPLA algorithm. Forcing estimation results are given in section 6.3. We conclude the paper in section 7.
@@ 2.1 Statistical Finite Elements for Linear PDEs
$\mathcal{A}_{\theta}(u_{h},\phi_{j})=\sum_{i=1}^{n_{u}}\hat{u}_{i}\mathcal{A}_{\theta}(\phi_{i},\phi_{j})=\left\langle f,\phi_{j}\right\rangle+\left\langle\epsilon,\phi_{j}\right\rangle,\quad\forall j\in\{1,\dots,n_{u}\},$ (8)
@@
Provided the bilinear form $\mathcal{A}_{\theta}$ is coercive on $V$ the problem is well-posed, the stiffness matrix $\mathbf{A}_{\theta}$ will be positive definite, implying its inverse exists (see e.g. [23][Remark 2.20]). For setting the homogeneous Dirichlet boundary conditions, we set the rows corresponding to boundary nodes to the row of the identity matrix, and the forcing vector to zero. This can be adapted for inhomogeneous boundary conditions via boundary condition lifting (see e.g. [30][Page 8]), which adapts the entries of the right-hand side vector to the non-zero boundary condition. We choose to constrain the covariance function $k$ to be zero on the boundary, so that boundary conditions are imposed exactly. To achieve this, we form an approximation to the GP, as seen in [44], which has the benefit in our case of being constrained to zero covariance on the boundary. The eigenfunctions of the Laplacian operator (essentially the vibrational modes of the domain) form the basis, which are weighted based on the spectrum of the covariance function. This approximation is given by
@@
where the function $g_{l}(x)=\sum_{l=1}^{n_{u}}\tilde{g}_{li}\phi_{i}(x)$ is the $L^{2}$-normalised FEM approximation of the $l^{\text{th}}$ eigenfunction of the Laplacian operator, $\lambda_{l}$ the corresponding eigenvalue, and $m$ the rank of the approximation. The weighting function $S(\cdot)$ is the spectral density of the covariance function $k$. Note this requires an isotropic covariance function, i.e. such that we can write $k(x,x^{\prime})=k(\|x-x^{\prime}\|)$. This is detailed in Appendix F.
@@ 2.2 Statistical Finite Elements via Langevin Dynamics
where $\left\{\bm{\zeta}_{k}\right\}_{k\in\mathbb{N}}$ is a sequence of i.i.d. standard Gaussian random variables, and $\gamma>0$ controls the degree of time-discretisation. This discretisation induces a bias, the degree of which is determined by the step-size $\gamma$. For some applications this bias may be unacceptable, and can be corrected by a Metropolis-Hastings style accept/reject step for each iteration of the algorithm (this is known as Metropolis adjusted Langevin algorithm (MALA) [27]) under the condition that the unnormalised density can be evaluated directly, not just the score as for the unadjusted Langevin algorithm (ULA). ULA does not perform this accept/reject stage and is therefore more computationally efficient than MALA, at the cost of introducing discretization bias [12, 13, 20].
@@
Preconditioning can improve the stability and convergence of the sampling scheme, whilst leaving the stationary measure invariant [4]. Section 3.1 shows how preconditioning is applied for posterior sampling.
@@ 2.3 Interacting Particle Langevin Dynamics
Whilst sections below will detail the IPLA algorithm for forcing estimation and solving the inverse problem, we provide a brief overview here. IPLA simultaneously evolves a set of particles with Langevin dynamics, which are used to estimate the gradient of the marginal likelihood with respect to parameters. A system of $N$ particles, $\{\mathbf{u}_{t}^{(n)}\}_{n=1}^{N}$ are simulated from the Langevin dynamics described in (15), with independent Brownian motions $(\mathbf{B}_{t}^{(n)})_{t\geq 0}$ for $n=1,\ldots,N$. These particles and parameters are jointly performing gradient flow on the negative joint log likelihood where the parameter process is driven with a scaled Brownian motion $\sqrt{{2}/{N}}\mathrm{d}\mathbf{B}_{t}^{(0)}$. This noise term, which is the main difference from the methods introduced in [34], allows for non-asymptotic analysis, and in [3], it is shown that given Lipschitz and strong convexity assumptions on the potential, the resulting parameter estimate will converge to the true MMLE, and the rate of convergence is derived in terms of the Lipschitz and strong convexity constants, and time-discretization step-size. The method has also shown to be adaptable to the nondifferentiable and underdamped cases, see, e.g. [21, 38]. Sections 3.1 and 3.2 give the resulting interacting particle systems for estimation of the forcing function and solving the inverse problem, respectively.
3 Interacting Particle Langevin Dynamics for Linear Poisson PDE
One example of an elliptic PDE is the Poisson equation, which describes processes such as steady-state heat diffusion. The strong form of the linear Poisson equation with homogeneous diffusivity coefficient $\varkappa$, is given by
$-\nabla\cdot(\varkappa\nabla u(x))$ $=f+\epsilon,\quad x\in\Omega$ (17)
$u(x)$ $=0,\quad x\in\partial\Omega.$ (18)
In our case we use the exponential mapping to reparameterise the strong form with diffusivity $\varkappa$, by $\theta=\log\varkappa$ to ensure non-negativity. The bilinear form associated with this strong form, $\mathcal{A}_{\theta}(u,v)=e^{\theta}\left\langle\nabla u,\nabla v\right\rangle$, produces the stiffness matrix $\left[\mathbf{A}_{\theta}\right]_{ij}=e^{\theta}\left\langle\nabla\phi_{i},\nabla\phi_{j}\right\rangle$. Since the bilinear form is symmetric, the stiffness matrix is symmetric positive definite, and the linear system, $\mathbf{A}_{\theta}\mathbf{u}=\mathbf{b}$, will have a unique solution.
The joint likelihood model, $p(\mathbf{y},\mathbf{u}|\mathbf{b},\theta)$, factorises as the product of the following Gaussian distributions
$p(\mathbf{u}|\theta,\mathbf{b})=\mathcal{N}\left(\mathbf{A}_{\theta}^{-1}\mathbf{b},\mathbf{A}_{\theta}^{-1}\mathbf{G}\mathbf{A}_{\theta}^{-\top}\right),\quad p(\mathbf{y}|\mathbf{u})=\mathcal{N}\left(\mathbf{H}\mathbf{u},\mathbf{R}\right),$ (19)
which defines the potential $\Phi_{\theta}^{\mathbf{y}}(\mathbf{u},\mathbf{b})$, which can be used for posterior sampling or for MMLE of either $\mathbf{b}$, $\theta$ using IPLA.
To further enable MMAP estimation of the forcing $\mathbf{b}$, and the parameter $\theta$, we require prior distributions over these quantities. We will assume independence $p(\mathbf{b},\theta)=p(\mathbf{b})p(\theta)$, as we will be estimating these individually. The discretised prior distribution over the forcing coefficients, $\mathbf{b}$, can be induced from a GP prior over the forcing function. Given $f\sim\mathcal{GP}(\mu_{f},k_{f})$, we have $\left[\bm{\mu}\right]_{i}=\langle\phi_{i},\mu_{f}\rangle$, $\left[\bm{\Sigma}\right]_{ij}=\langle\phi_{i},\langle k_{f}(\cdot,\cdot),\phi_{j}\rangle\rangle$. For the parameter $\theta$ we assume a Gaussian prior, which is equivalent to a log-normal prior over the diffusivity, but with the added constraint of non-negativity when performing gradient update steps. The prior distributions are as follows
$p(\mathbf{b})=\mathcal{N}\left(\bm{\mu},\bm{\Sigma}\right),\quad p(\theta)=\mathcal{N}\left(\mu_{\theta},\sigma^{2}_{\theta}\right).$ (20)
We have now defined each component of the full joint probabilistic model $p(\mathbf{y},\mathbf{z},\mathbf{u},\mathbf{b},\theta)=p(\mathbf{y}|\mathbf{u})p(\mathbf{u}|\mathbf{b},\theta)p(\mathbf{b})p(\theta)$, which we use to define the joint potential for the MMAP estimation, using the potential in (15) for posterior sampling, and the prior distributions $p(\mathbf{b}),p(\theta)$:
$\Psi_{\theta}^{y}(\mathbf{u},\mathbf{b}):=\Phi_{\theta}^{y}(\mathbf{u},\mathbf{b})-\log p(\mathbf{b})-\log p(\theta).$ (21)
Using the forcing estimation as an example to illustrate the standard IPLA, we present the interacting Langevin SDE for estimation of $\mathbf{b}$, with $\theta$ fixed between iterations. For a system of $N$ particles, $\{\mathbf{u}_{t}^{(n)}\}_{n=1}^{N}$, the interacting Langevin SDE has the following form
$\mathrm{d}\mathbf{b}_{t}$ $=-\frac{1}{N}\sum_{n=1}^{N}\nabla_{\mathbf{b}}\Psi_{\theta}^{y}(\mathbf{u}_{t}^{(n)},\mathbf{b}_{t})\mathrm{d}t+\sqrt{\frac{2}{N}}\,\mathrm{d}\mathbf{B}_{t}^{(0)}$ (22)